OpenStudy (anonymous):

Simplify completely 3x^2 + 21x -54/x^2+3x-54

OpenStudy (cornzyblack):

well 9x^2 + 21x - (54/x^2) + 3x - 54 = collecting like terms, 9x^2 - (54/x^2) + 3x + 21x - 54= (9x^4 - 54)/x^2 + 24x - 54= 3([(3x^4 - 18)/x^2] + 8x - 18)